OpenStudy (anonymous):

Solve by completing the square x2 − 9x + 3 = 0

OpenStudy (anonymous):

x2 - 9x + (9/2)^2 + 3 = (9/2)^2 (x - 4.5)^2 = 17.25 x-4.5 = +- 4.15 x = 8.65 or x = 0.35